OpenStudy (erniellyd28):

recall that two angles are complementary if the sum of their measures is 90 degrees. find the measures of two complementary angles if one angle is seventeen times the other angle. (i found the smaller angle but not sure how to find the larger angle)

OpenStudy (hafsahh):

x+y=90 y=17x x+17x=90 18x=90 90/18=5 x=5 y=17*5 y=85 the angles are 5 and 85

OpenStudy (erniellyd28):

thank you very much for your help :-)